# A Consistent Cosmic Shear Analysis in Harmonic and Real Space

### v1

@@ 3.1.3 Accounting for Scale Cuts
The pseudo-$C_{\ell}$ measurements in equations (15) and (16) are inconsistent with traditional TPCFs measurements, and the two two-point estimators are weakly correlated (e.g., Hamana et al., 2020; Doux et al., 2021). This is because TPCFs are measured over a limited $\theta$ range in real space, but this is not explicitly modeled in the pseudo-$C_{\ell}$ measurements. Therefore, the two measurements effectively use different sets of information that are probed by different $\ell$ ranges. As shown in Singh (2021), to make a consistent two-point estimator in harmonic space with the TPCFs in real space, we need to account for the finite range of $\theta$ used in real space. Applying scale cuts in real space corresponds, in practice, to multiplying the TPCFs by a $\theta$-dependent function denoted as $\nu(\theta)$. Mathematically, this operation is represented as $\hat{\xi}_{\pm}(\theta)=\xi_{\pm}(\theta)\nu(\theta)$, where the scale cut $\nu(\theta)$ is typically a step function defined over a finite range of scales.
@@ 3.1.4 Binning in $\ell$
Figure 1: In the upper panel, we compare three different $E$-mode auto power spectra for the lowest tomographic bin $(0.3,0.6]$ of the fiducial cosmology defined in Sec. 4.2: $f_{\mathrm{sky}}C^{EE}_{\ell}$ (blue) with normalization convention of the pseudo-$C_{\ell}$ that includes a factor of $f_{\mathrm{sky}}$, $D^{EE}_{\ell}$ (orange), and $F^{EE}_{\ell}$ (green). The $B$-mode leakage of the form $M^{-,sc}M^{-,sw}C^{EE}_{\ell}$ in equation (21) is plotted as a dotted red line. We find that $B$-mode leakage contributes $\sim 5\%$ to the $E$-mode consistent pseudo-$C_{\ell}$. In the lower panel, we plot the ratios between the power spectrum and pseudo-$C_{\ell}$ (orange) and between pseudo-$C_{\ell}$ and consistent pseudo-$C_{\ell}$ (green). We find that information is lost on large scales when measuring the power spectrum on a weighted map, and power is recovered on small scales. Power is further lost on small scales when we model the impact of the scale cut used in real space in harmonic space.
@@ 3.2 Correlation Function
Traditional TPCFs estimators return TPCFs that are average weighted by the number of pairs at each angular bin. However, the number of galaxies in each angular bin can be noisy with many bins, and the TPCFs estimator defined in equation (3.2) does not fully remove the effects of the survey window. This is due to the effects of separately binning the numerator and denominator in equation (3.2). This way, the window effect is accounted for imperfectly, which leads to residual percent-level window effects when using traditional TPCFs estimators (see appendix D of Singh et al. 2020 for demonstration of residual percent-level window effects on correlation functions). Rather than trying to use the above estimator and also identify and apply those additional window corrections, we take the alternate approach with a modified estimator from Singh (2021) that is consistent with the pseudo-$C_{\ell}$ defined in equations (15) and (16), and is motivated by the weighting scheme in equation (14).
@@ 3.3 Going from the TPCFs to the Power Spectrum
Then, using equations (26) and (27), we can reconstruct the underlying cosmic shear power spectrum.
@@ 3.4 Going from the Power Spectrum to the TPCFs
Similar to how we manually convolved the pseudo-$C_{\ell}$ to account for the scale cut used in real space, we can simply convolve the measured TPCFs with $b_{\theta\theta^{\prime}}$ to impose the scale cut used in harmonic space before running the inference chain.
@@ 4.3 Power Spectra From Pseudo Power Spectra
We use the pseudo-$C_{\ell}$ method described in Sec. 3.1 to reconstruct the tomographic cosmic shear power spectra from the HSC Y1 shape mock catalogs. We measure the power spectra and mode coupling matrices to larger $\ell$ than we wish to use for our analysis to correctly account for mode coupling between smaller scale modes and to account for the $E/B$-mode leakage. The shear maps of each tomographic bin of the mocks were constructed using Healpy (Górski et al., 2005) following equation (12) with a resolution parameter $N_{\text{side}}=4096$, corresponding to a pixel size of $\sim 0.86^{\prime}$. We measure the tomographic full-sky pseudo power spectra up to $\ell_{\text{max}}=8000$, and the same $\ell_{\text{max}}$ was used to generate the coupling matrices $M^{sc}_{\ell\ell^{\prime}}$ and $M^{sw}_{\ell\ell^{\prime}}$ in equation (17). Before reconstructing the power spectra from the pseudo power spectra, we correct for the pixel window function by dividing the power spectra in equations (15) and (16) by the angular power spectrum of the pixel window function; $C^{(\Pi)}_{\ell}\simeq\operatorname{sinc}(\ell\Delta\theta/2\pi)$ is the pixel window function of Healpix (Jeong et al., 2014), where $\Delta\theta$ is the angular resolution. We then convolve the pseudo power spectra with a coupling matrix derived from the scale cut (equations (3.1.3) and (20)) to estimate the consistent pseudo-$C_{\ell}$ and bin each spectrum and coupling matrices as described by the binning operator in equation (24). Equations (26) and (27) are then used to reconstruct the $EE$ and $BB$ spectra for each tomographic bin pair.
@@ 4.4 Power Spectra From TPCFs
Once we have our TPCFs, we multiply by the survey window correlation function to make it consistent with the power spectra measurements, resulting in measurements as described in equation (31). After binning the measurements and the inverse spherical transform operator, equations (34) and (35) are used to get the pseudo-$D_{\ell}$, in which we reconstruct the power spectra using equations (26) and (27). We reconstruct the $EE$ and $BB$ spectra in the same 20 linearly equal $\ell$ bins defined in 4.3, resulting in 70 data points for the data vector for each mock.
@@ 4.5 Covariance Matrix
The lensing covariance matrix can be decomposed into three terms: the Gaussian, connected non-Gaussian, and the super-sample covariance (Barreira et al. 2018). The Gaussian covariance term of the cosmic shear power spectrum can be further decomposed into three terms: the auto-term of cosmic shear, the cross-term of cosmic shear and shape noise spectra, and the auto-term of the shape noise spectra as
@@
Figure 3: Comparison of the reconstructed tomographic cosmic shear power spectra of $EE$ mode from pseudo power spectra and TPCFs measurements with error bars in the lower left panel. The colored points show the average reconstructed power spectrum from 571 mock realizations from pseudo power spectra (blue dots) and two-point correlation functions (orange dots) for the lowest redshift bin, as indicated by the $1\times 1$ label. The analytical forward-modeled prediction (solid black line) is also shown for comparison. The scales outside of $1800400<\ell<$ 1800 (shaded regions) are excluded from the inference. The other panels show the ratio of each reconstruction to the forward-modeled theory, with error bars. The error bars are $1\sigma$ uncertainties in the ratio estimated using mock catalogs. Reconstructions of power spectra from $i$Master give consistent results to well within $1\%$ and they give unbiased results to within $2\%$. Black dotted lines represent the $2\%$ bias region. The scales outside of the region $1800400<\ell<$ 1800 (shaded regions) are excluded from the cosmological analysis.
@@ 5.1 Cosmic Shear Power Spectrum Reconstruction
where $q\in\{\text{PS},\text{TPCF}\to\text{PS}\}$ represents the direction of the analysis and $\Sigma$ is the covariance matrix estimated from the mocks. The two integrated S/N of the reconstructions are in agreement within ±$\pm 1\%$. The S/N of the reconstructed power spectrum from the consistent pseudo-$C_{\ell}$ in individual redshift bins are 4.86, 9.32, 11.68, and 10.09 from the lowest to highest redshift bins, respectively. The S/N in these redshift bins are consistent with those of the HSC Y1 power spectrum and correlation function analyses to within $20\%$. The S/N from mocks is expected to be less than that of real data, as real data contain noise, and the noise contributes to S/N. To quantify the covariance of two reconstructions, we compute the correlation coefficient as
@@
with $x=\hat{C}_{\ell}$. The subscripts $R$ and $F$ stand for the reconstruction from TPCFs (real space) and consistent pseudo-$C_{\ell}$ ( harmonic space), respectively. We find that the correlation coefficient range is $r(\hat{C}_{\ell})\in[0.91,1.0]$, thus the two reconstructions have the same information and are consistent. For a detailed review, refer to Appendix A.
@@ 5.2 Likelihood Inference
Figure 5: Left panel: scatter plot showing the inferred $S_{8}$ values from consistent pseudo-$C_{\ell}$ analysis in the horizontal axis and from two-point correlation functions analysis in the vertical axis. The points are from minimizing $\chi^{2}$ for 571 mock catalogs. The red $+$ denotes the value of $S_{8}$ adopted in generating the mock catalogs. The inferred $S_{8}$ values between the two independent analyses are very consistent, with the correlation coefficient $r=0.994$. Right panel: we show the distribution of $S_{8}$ from the mocks for each axes. The dotted vertical lines represent the mean of each distribution, with the red dotted vertical line denoting the fiducial $S_{8}$ value. The two distributions have very similar means, with $\Delta\langle S_{8}\rangle=0.0004$.
@@ 6 Conclusion
In this paper, we presented the first practical application of the $i$Master algorithm (Singh, 2021) to mock catalogs with a realistic cosmological density field and a realistically complex survey window (the HSC Y1 mock shape catalogs). In doing so, we reviewed the formalism for the measurement of cosmic shear two-point functions and how to correctly account for survey window and scale cuts in harmonic and real space to obtain a consistent power spectrum.
In Sec. 3, we discussed the impact of the survey window and scale cut on the various estimators. Applying the survey window and scale cuts in real space generally involves multiplying the underlying maps by some weights, which are equivalent to convolutions in harmonic space. Since measurements of the power spectrum with non-uniform coverage of the sky do not account for the scale cut used in real space, we manually convolved the pseudo-$C_{\ell}$ to obtain the consistent pseudo-$C_{\ell}$. We showed that this greatly affects the pseudo-$C_{\ell}$ measurements in all scales. We also showed that leakage of the form $M^{-,sc}M^{-,sw}C^{EE}_{\ell}$, equivalent to $B$-mode leakage of $F^{EE}_{\ell}$, contributes $5$ per cent to the $F^{EE}_{\ell}$ measurements. We also derived a new two-point correlation function (TPCF) estimator that is consistent with the process of estimating the pseudo-$C_{\ell}$. In particular, the new estimator normalizes the measurements by the expected number of galaxy pairs instead observed number of galaxy pairs, which is traditionally done in the literature. This new procedure changes the two-point correlation function measurements by $\sim$2 per cent and allows for self-consistent modeling in the two spaces.
@@
In Sec. 5, we showed the result of applying the $i$Master algorithm using the HSC Y1 shape mock catalog. Repeating the scale cuts used in both harmonic and real space in the actual HSC Y1 analysis, $i$Master algorithm was able to reconstruct the cosmic shear power spectrum from $F^{EE/BB}_{\ell}$ and $\xi_{\pm}$ that agree within ±$\pm 1$ per cent with each other. We also showed that the two integrated S/N of the reconstructions are consistent within ±$\pm 1$ per cent. We compared the inferred $S_{8}$ values using likelihood inference with the reconstructed cosmic shear power spectrum. We found that the two inferred $S_{8}$ values from 571 mocks to be strongly correlated with a correlation coefficient of 0.994.

### v2

@@ 3.1.3 Accounting for Scale Cuts
The pseudo-$C_{\ell}$ measurements in equations (15) and (16) are not matching with traditional TPCFs measurements, and the two two-point estimators are weakly correlated (e.g., Hamana et al., 2020; Doux et al., 2021). This is because TPCFs are measured over a limited $\theta$ range in real space, but this is not explicitly modeled in the pseudo-$C_{\ell}$ measurements. Therefore, the two measurements effectively use different sets of information that are probed by different $\ell$ ranges. As shown in Singh (2021), to make a matching two-point estimator in harmonic space with the TPCFs in real space, we need to account for the finite range of $\theta$ used in real space. Applying scale cuts in real space corresponds, in practice, to multiplying the TPCFs by a $\theta$-dependent function denoted as $\nu(\theta)$. Mathematically, this operation is represented as $\hat{\xi}_{\pm}(\theta)=\xi_{\pm}(\theta)\nu(\theta)$, where the scale cut $\nu(\theta)$ is typically a step function defined over a finite range of scales.
@@ 3.1.4 Binning in $\ell$
Figure 1: In the upper panel, we compare three different $E$-mode auto power spectra for the lowest tomographic bin $(0.3,0.6]$ of the fiducial cosmology defined in Sec. 4.2, using the HSC Y1 mask: $f_{\mathrm{sky}}C^{EE}_{\ell}$ (blue) with normalization convention of the pseudo-$C_{\ell}$ that includes a factor of $f_{\mathrm{sky}}$, $D^{EE}_{\ell}$ (orange), and $F^{EE}_{\ell}$ (green). The $B$-mode leakage of the form $M^{-,sc}M^{-,sw}C^{EE}_{\ell}$ in equation (21) is plotted as a dotted red line. We find that $B$-mode leakage contributes $\sim 5\%$ to the $E$-mode consistent pseudo-$C_{\ell}$. In the lower panel, we plot the ratios between the power spectrum and pseudo-$C_{\ell}$ (orange) and between pseudo-$C_{\ell}$ and consistent pseudo-$C_{\ell}$ (green). We find that information is lost on large scales when measuring the power spectrum on a weighted map, and power is recovered on small scales. Power is further lost on small scales when we model the impact of the scale cut used in real space in harmonic space.
@@ 3.2 Correlation Function
Traditional TPCFs estimators return TPCFs that are average weighted by the number of pairs at each angular bin. However, the number of galaxies in each angular bin can be noisy with many bins, and the TPCFs estimator defined in equation (3.2) does not fully remove the effects of the survey window. This is due to the effects of separately binning the numerator and denominator in equation (3.2). This way, the window effect is accounted for imperfectly, which leads to residual percent-level window effects when using traditional TPCFs estimators (see appendix D of Singh et al. 2020 for demonstration of residual percent-level window effects on correlation functions). Rather than trying to use the above estimator and also identify and apply those additional window corrections, we take the alternate approach with a modified estimator from Singh (2021) that is matching with the pseudo-$C_{\ell}$ defined in equations (15) and (16), and is motivated by the weighting scheme in equation (14). Using an estimator that explicitly accounts for the survey window ensures matching results with harmonic space analyses.
@@ 3.3 Going from the TPCFs to the Power Spectrum
Then, using equations (26) and (27), we can reconstruct the underlying cosmic shear power spectrum. By accounting for the survey window and scale cuts, our method ensures that the reconstructed power spectrum is matching, allowing for consistent cosmic shear analysis.
@@ 3.4 Going from the Power Spectrum to the TPCFs
Similar to how we manually convolved the pseudo-$C_{\ell}$ to account for the scale cut used in real space, we can simply convolve the measured TPCFs with $b_{\theta\theta^{\prime}}$ to impose the scale cut used in harmonic space before running the inference chain. By properly handling scale-dependent effects, we ensure that real space and harmonic space measurements provide matching constraints on cosmological parameters.
@@ 4.3 Power Spectra From Pseudo Power Spectra
We use the pseudo-$C_{\ell}$ method described in Sec. 3.1 to reconstruct the tomographic cosmic shear power spectra from the HSC Y1 shape mock catalogs. We measure the power spectra and mode coupling matrices to larger $\ell$ than we wish to use for our analysis to correctly account for mode coupling between smaller scale modes and to account for the $E/B$-mode leakage. The shear maps of each tomographic bin of the mocks were constructed using Healpy (Górski et al., 2005) following equation (12) with a resolution parameter $N_{\text{side}}=4096$, corresponding to a pixel size of $\sim 0.86^{\prime}$. We measure the tomographic full-sky pseudo power spectra up to $\ell_{\text{max}}=8000$, and the same $\ell_{\text{max}}$ was used to generate the coupling matrices $M^{\nu}_{\ell\ell^{\prime}}$ and $M^{w}_{\ell\ell^{\prime}}$ in equation (17). Before reconstructing the power spectra from the pseudo power spectra, we correct for the pixel window function by dividing the power spectra in equations (15) and (16) by the angular power spectrum of the pixel window function; $C^{(\Pi)}_{\ell}\simeq\operatorname{sinc}(\ell\Delta\theta/2\pi)$ is the pixel window function of Healpix (Jeong et al., 2014), where $\Delta\theta$ is the angular resolution. We then convolve the pseudo power spectra with a coupling matrix derived from the scale cut (equations (3.1.3) and (20)) to estimate the consistent pseudo-$C_{\ell}$ and bin each spectrum and coupling matrices as described by the binning operator in equation (24). Equations (26) and (27) are then used to reconstruct the $EE$ and $BB$ spectra for each tomographic bin pair.
@@ 4.4 Power Spectra From TPCFs
Once we have our TPCFs, we multiply by the survey window correlation function to match with the power spectra measurements, resulting in measurements as described in equation (31). After binning the measurements and the inverse spherical transform operator, equations (34) and (35) are used to get the pseudo-$D_{\ell}$, in which we reconstruct the power spectra using equations (26) and (27). We reconstruct the $EE$ and $BB$ spectra in the same 20 linearly equal $\ell$ bins defined in 4.3, resulting in 70 data points for the data vector for each mock.
@@ 4.5 Covariance Matrix
The lensing covariance matrix can be decomposed into three terms: the Gaussian, connected non-Gaussian, and the super-sample covariance (Barreira et al. 2018). The non-Gaussian terms are not modeled as the covariance matrix estimation is done on Gaussian mocks. The Gaussian covariance term of the cosmic shear power spectrum can be further decomposed into three terms: the auto-term of cosmic shear, the cross-term of cosmic shear and shape noise spectra, and the auto-term of the shape noise spectra as
@@
Figure 3: Comparison of the reconstructed tomographic cosmic shear power spectra of $EE$ mode from pseudo power spectra and TPCFs measurements with error bars in the lower left panel. The colored points show the average reconstructed power spectrum from 571 mock realizations from pseudo power spectra (blue dots) and two-point correlation functions (orange dots) for the lowest redshift bin, as indicated by the $1\times 1$ label. The analytical forward-modeled prediction (solid black line) is also shown for comparison. The scales outside of $1800400<\ell<$ 1800 (shaded regions) are excluded from the inference. The other panels show the ratio of each reconstruction to the forward-modeled theory, with error bars. The error bars are $1\sigma$ uncertainties in the ratio estimated using mock catalogs. Reconstructions of power spectra from $i$Master give matching results to well within $1\%$ and they give unbiased results to within $2\%$. Black dotted lines represent the $2\%$ bias region. The scales outside of the region $1800400<\ell<$ 1800 (shaded regions) are excluded from the cosmological analysis.
@@ 5.1 Cosmic Shear Power Spectrum Reconstruction
where $q\in\{\text{PS},\text{TPCF}\to\text{PS}\}$ represents the direction of the analysis and $\Sigma$ is the covariance matrix estimated from the mocks. The two integrated S/N of the reconstructions are in agreement within ±$\pm 1\%$. The S/N of the reconstructed power spectrum from the consistent pseudo-$C_{\ell}$ in individual redshift bins are 4.86, 9.32, 11.68, and 10.09 from the lowest to highest redshift bins, respectively. The S/N in these redshift bins are consistent with those of the HSC Y1 power spectrum and correlation function analyses to within $20\%$. The S/N from mocks is expected to be less than that of real data, as real data contains noise that artificially biases the measured signal to noise ratio to higher values. To quantify the covariance of two reconstructions, we compute the correlation coefficient as
@@
with $x=\hat{C}_{\ell}$. The subscripts $R$ and $F$ stand for the reconstruction from TPCFs (real space) and consistent pseudo-$C_{\ell}$ (harmonic space), respectively. We find that the correlation coefficient range is $r(\hat{C}_{\ell})\in[0.91,1.0]$, thus the two reconstructions have the same information and are matching. For a detailed review, refer to Appendix A.
@@ 5.2 Likelihood Inference
Figure 5: Left panel: scatter plot showing the inferred $S_{8}$ values from consistent pseudo-$C_{\ell}$ analysis in the horizontal axis and from two-point correlation functions analysis in the vertical axis. The points are from minimizing $\chi^{2}$ for 571 mock catalogs. The red $+$ denotes the value of $S_{8}$ adopted in generating the mock catalogs. The inferred $S_{8}$ values between the two independent analyses are very consistent, with the correlation coefficient $r=0.994$. While the comparison is conceptually similar to Figure 17 of (Hamana et al., 2020), which highlights the inconsistency between different summary statistics, we emphasize that their analysis was performed on noisy mock data, whereas our analysis uses noiseless simulations. As such, the results are not directly comparable. Right panel: we show the distribution of $S_{8}$ from the mocks for each axes. The dotted vertical lines represent the mean of each distribution, with the red dotted vertical line denoting the fiducial $S_{8}$ value. The two distributions have very similar means, with $\Delta\langle S_{8}\rangle=0.0004$.
@@ 6 Conclusion
In this paper, we presented the first practical application of the $i$Master algorithm (Singh, 2021) to mock catalogs with a realistic cosmological density field and a realistically complex survey window (the HSC Y1 mock shape catalogs). In doing so, we reviewed the formalism for the measurement of cosmic shear two-point functions and how to correctly account for survey window and scale cuts in harmonic and real space to obtain a matching power spectrum.
In Sec. 3, we discussed the impact of the survey window and scale cut on the various estimators. Applying the survey window and scale cuts in real space generally involves multiplying the underlying maps by some weights, which are equivalent to convolutions in harmonic space. Since measurements of the power spectrum with non-uniform coverage of the sky do not account for the scale cut used in real space, we manually convolved the pseudo-$C_{\ell}$ to obtain the consistent pseudo-$C_{\ell}$. We showed that this greatly affects the pseudo-$C_{\ell}$ measurements in all scales. We also showed that leakage of the form $M^{-,sc}M^{-,sw}C^{EE}_{\ell}$, equivalent to $B$-mode leakage of $F^{EE}_{\ell}$, contributes $5$ per cent to the $F^{EE}_{\ell}$ measurements. We also derived a new two-point correlation function (TPCF) estimator that is matching with the process of estimating the pseudo-$C_{\ell}$. In particular, the new estimator normalizes the measurements by the expected number of galaxy pairs instead observed number of galaxy pairs, which is traditionally done in the literature. This new procedure changes the two-point correlation function measurements by $\sim$2 per cent and allows for self-consistent modeling in the two spaces.
@@
In Sec. 5, we showed the result of applying the $i$Master algorithm using the HSC Y1 shape mock catalog. Repeating the scale cuts used in both harmonic and real space in the actual HSC Y1 analysis, $i$Master algorithm was able to reconstruct the cosmic shear power spectrum from $F^{EE/BB}_{\ell}$ and $\xi_{\pm}$ that match within ±$\pm 1$ per cent with each other. We also showed that the two integrated S/N of the reconstructions are matching within ±$\pm 1$ per cent. We compared the inferred $S_{8}$ values using likelihood inference with the reconstructed cosmic shear power spectrum. We found that the two inferred $S_{8}$ values from 571 mocks to be strongly correlated with a correlation coefficient of 0.994.